OpenStudy (anonymous):

f(x)=x/2-x. Find f(1/x)

OpenStudy (anonymous):

Just plug in "1/x" in everywhere "x" shows up: \[f(x)=f(1/x)=(1/x)/[2-(1/x)]\]

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

what else

OpenStudy (anonymous):

That reduces down to \[1/x(2-1/x) = 1/(2x-1)\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

multiple both sides by x?

OpenStudy (anonymous):

Starting with the original \[(1/x)/[2-(1/x)]\] it's just elementary math: instead of diving by [2-(1/x)], multiply by its reciprocal 1/[2-(1/x)]. So it will look like this: \[(1/x)/[2-(1/x)] = (1/x) * 1/[2-(1/x)] = 1/[x(2-(1/x))] = 1/(2x-x/x) = 1/(2x-1)\]

OpenStudy (anonymous):

Oops, did that get cut off on the right side? The end is supposed to be 1/(2x-x/x) = 1/(2x-1)

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

You're welcome. I hope it made sense! When you have fractions in both a numerator and denominator, multiplying by the denominator's reciprocal instead of leaving it as a division is a good way to simplify things.

OpenStudy (anonymous):

:)